OpenStudy (cp9454):

Between two numbers a & b first n1 geometric means are inserted. Then between the same number n2 geometric means are inserted. find the ratio of product of geometric means.

OpenStudy (anonymous):

first find the common ratio in each case...

OpenStudy (cp9454):

i think commen ratio will be same in both the cases, only number of term is changing.

OpenStudy (anonymous):

no, common ratio will be differnt

OpenStudy (cp9454):

ok then, what

OpenStudy (anonymous):

in the first case first term is a and b is (n1+2)th term and no of terms is (n1+2) in the second case first term is a and b is (n2+2)th term and no of terms is (n2+2)

OpenStudy (anonymous):

let r1 and r2 be the common ratio in the above two cases..

OpenStudy (anonymous):

then b = a( (r1)^(n1+1)) and b = a ((r2)^(n2+1)) and so on..